OpenStudy (anonymous):

solve 12 sin x cos x using the double-angle formula

OpenStudy (jdoe0001):

\(\bf sin(2\theta)={\color{blue}{ 2sin(\theta)cos(\theta)}}\\ \quad \\ \quad \\ 12sin(x)cos(x)\implies 6\cdot {\color{blue}{ 2sin(x)cos(x)}}\)